# Berkeley, CA 94720, USA

### v2

@@ 6 Conclusions and Further Directions
In this paper, we built on work of Bak and Gustavsson Bak:2017jwd , combined with results of Chang and Dowker Dowker:2009in on the Ray-Singer torsion of spherical orbifolds, to compute the Witten index of various systems obtained from the 6d (2,0)-theory associated to gauge group SU$(2)$. We compactified the 6d theory on 5-manifolds $X_{5}$ constructed from spherical orbifolds with a SUSY preserving twist, and with possible insertions of surface operators playing the role of external sources. A connection between the values of the Ray-Singer torsion and the characters of representations of the orbifold groups, in terms of which the Witten index is expressed, ensures that the result is always an integer.
@@
Our main example is $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, where $\text{Dic}_{k}$ is the dicyclic group (of order $4k$), and in this case the results for the Witten index also allow us to postulate conjectures regarding fusion relations among external string-like sources, represented by timelike surface operators. We also discussed $S^{5}/\mathbb{Z}_{k}$ and $S^{3}/\mathbb{Z}_{k}\times S^{2}$, where the Witten index is zero, because of fermionic zero modes. Extensions to $S^{5}/\Gamma$ and $S^{3}/\Gamma\times S^{2}$ with other (non-abelian) finite orbifold groups $\Gamma$ are in progress WIP .
@@
Our strategy, which follows Bak:2017jwd , was to reduce the problem to a partition function of topologically twisted 5d SYM on $X_{5}$, containing gauge fields, bosonic and Grassmann odd 0-forms, 1-forms and 2-forms which participate in a 5d “Fermionic Chern Simons” term in the action. Supersymmetry ensures that the partition function localizes on gauge field configurations that are flat connections, and a careful treatment of fermionic zero modes shows that only irreducible non-abelian flat connections contribute. The simplest non-trivial case is then $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, since for $\mathbb{Z}_{k}$ quotient, all representations are abelian. Using expressions developed in Bak:2017jwd for the one-loop contribution in terms of the Ray-Singer analytic torsion, we recast the partition function as a sum over non-abelian representations of Dick torsions (associated with a flat connection) times products of characters of group elements (38). We simplified the torsion results of Dowker:2009in for $S^{3}/\text{Dic}_{k}$ to a compact expression (62), and we used it to confirm that the Witten index is indeed given by an integer. Recurrence relations among the Witten indices for various numbers of timelike surface operator insertions lead to conjectural fusion rules for the external string sources associated with those operators. These relations are expressed as combinations of Chebyshev polynomials.
Moreover, the results obtained in this work provide a physical explanation for the existence of “nice” closed form expressions for the Ray-Singer analytic torsion for $S^{3}/\text{Dic}_{k}$. In our framework, the simplicity of these expressions is directly linked to the fact that physics requires the Witten index to be an integer for any number of external surface operator insertions. Note that the Ray-Singer torsion itself is a sine of a rational multiple of $\pi$, which is generally an irrational number. This situation is perhaps reminiscent of the Verlinde formula Verlinde:1988sn and its application to 3d Chern-Simons theory Witten:1988hf , where integer fusion coefficients $N_{ijk}$ are expressed as sums over representations and matrix elements of modular transformations that are sines of rational multiples of $\pi$. One could potentially rederive the formulas in this paper by representing $S^{3}/\text{Dic}_{k}$ as a $\mathbb{Z}_{2}$ quotient of the Hopf fibration, and reducing from 6d to 5d on the circle fiber of the Hopf fibration. This leads to a 5d SYM theory on ×$\mathbb{R}\times S^{2}\times\mathbf{D}^{2}$, where $\mathbf{D}^{2}$ is a $2$-disk, with the antipodal points of $S^{1}=\partial\mathbf{D}^{2}$ identified by complex conjugation on the fields. One can show that this leads to an effective Chern-Simon theory at level $(k-2)$ which can explain the appearance of the sines WIP .
@@
It would be interesting to understand operators with nontrivial matrix elements between states of space of ground states that we constructed, and it might provide a step towards a fundamental formulation of the $(2,0)$-theory, at least in certain low-energy limits. Looking ahead, our analysis may be naturally extended in several directions. In particular, one could study the smooth orbifolds $X_{5}=S^{5}/\Gamma$ for the non-abelian $\Gamma$’s that appear in (8). Such an extension is expected to involve similar techniques, namely the computation of the Ray-Singer torsion in the background of flat connections associated to representations of $\Gamma$, while potentially uncovering further insights into the structure of the underlying $(2,0)$ theory. Our results can also be examined in the context of 4d $\mathcal{N}=2$ $\text{Class}\;\mathcal{S}$ theories Gaiotto:2009we ; Gaiotto:2009hg , by first compactifying on $S^{2}$ (with a topological twist), leading to a gapped 4d theory. In that case, the Witten index gives information about the non-trivial topological nature of this gapped system.
@@
We have worked with an SU$(2)$ gauge group, because for SU$(N)$ with $N\geq 3$ the fermionic 0-forms will always have zero modes. This is because there are no irreducible representations of $\text{Dic}_{k}$ of dimension higher than $2$, and therefore for any representation $\mathbf{R}$ of $\text{Dic}_{k}$ there are non-trivial elements of the Lie algebra $\mathfrak{su}(N)$ (in which the 0-form takes values) that are invariant under the adjoint action of $\text{Dic}_{k}$ [embedded in SU$(N)$ via $\mathbf{R}$]. Moreover, one may consider more general surface operators Drukker:2020bes . It would also be interesting to find explicit operators that have non-trivial matrix elements between different ground states, and a natural candidate would be a surface operator wrapped on $S^{2}$. Such studies could reveal a richer structure in the spectrum of BPS states and their associated indices.
@@ Acknowledgements.
We wish to thank Vi Hong, Petr Hořava and Jesus Sanchez Jr. for helpful discussions. This work has been supported by the Berkeley Center for Theoretical Physics.

### v3

@@ 8 Conclusions and Further Directions
In this paper, we built on the work of Bak and Gustavsson Bak:2017jwd , combined with results of Chang and Dowker Dowker:2009in on the Ray-Singer torsion of spherical orbifolds, to compute the Witten index of various systems obtained from the 6d (2,0)-theory associated to gauge group SU(2)(2). We compactified the 6d theory on 5-manifolds $X_{5}$ constructed from spherical orbifolds with a SUSY preserving twist, and with possible insertions of surface operators playing the role of external sources. A connection between the values of the Ray-Singer torsion and the characters of representations of the orbifold groups, in terms of which the Witten index is expressed, ensures that the result is always an integer.
@@
Our main example is for the manifold $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, where $\text{Dic}_{k}$ is the dicyclic group (of order $4k$), and in this case the results for the Witten index also allow us to postulate conjectures regarding fusion relations among external string-like sources, represented by timelike surface operators. We also discussed $S^{5}/\mathbb{Z}_{k}$ and $S^{3}/\mathbb{Z}_{k}\times S^{2}$, where the Witten index is zero, because of fermionic zero modes. Extensions to $S^{5}/\Gamma$ and $S^{3}/\Gamma\times S^{2}$ with other (non-abelian) finite orbifold groups $\Gamma$ are possible future developments of this work.
@@
Our initial strategy, which follows Bak:2017jwd , was to reduce the problem to a partition function of topologically twisted 5d SYM on $X_{5}$, containing gauge fields, bosonic and Grassmann odd 0-forms, 1-forms and 2-forms which participate in a 5d “Fermionic Chern Simons” term in the action. Supersymmetry ensures that the partition function localizes on gauge field configurations that are flat connections, and a study of fermionic zero modes shows that only irreducible non-abelian flat connections contribute. The simplest nontrivial case is then $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, since for $\mathbb{Z}_{k}$ quotient, all representations are abelian. Using expressions developed in Bak:2017jwd for the one-loop contribution in terms of the Ray-Singer analytic torsion, we recast the partition function as a sum (42) over non-abelian representations of Dick torsions (associated with a flat connection) times products of characters of group elements. We simplified the torsion results of Dowker:2009in for $S^{3}/\text{Dic}_{k}$ to a compact expression (66), and we used it to confirm that the Witten index is indeed given by an integer. Recurrence relations among the Witten indices for various numbers of timelike surface operator insertions lead to conjectural fusion rules for the external string sources associated with those operators. These relations are expressed as combinations of Chebyshev polynomials.
Moreover, we were able to rederive the same expressions for the Witten indices from a dual perspective that led to 3d Chern-Simons theory with two different sectors, labeled by an SU(2)(2) element $\Omega$} that determines boundary conditions on fields, and gauge groups that depend on the parity of $k$. Thus, the results obtained in this work provide a physical explanation for the existence of “nice” closed form expressions for the Ray-Singer analytic torsion for $S^{3}/\text{Dic}_{k}$, and they follow from the dimensions of Hilbert spaces of Chern-Simons theory on $S^{2}$ with quark sources, as computed in Witten:1988hf . The fusion relations among surface operators then must follow from the basic fusion relations of Wilson loops in Chern-Simons theory, and can be expressed in terms of the $N_{ijk}$ fusion coefficients and Verlinde’s formula Verlinde:1988sn ; Witten:1988hf . This connection between Chern-Simons theory and analytic torsion on $S^{3}/\text{Dic}_{k}$ has, as far as we know, not been reported before in the literature.
It is interesting to reflect on how the expression (92), derived from analytic torsion in §5.2, and the derivation of the same formula from Chern-Simons theory in §6 are connected. The factor $\cos\left(\frac{2m\pi}{k}\right)$ in (92) arises from a certain character of $\text{Dic}_{k}$ in a representation labeled by $m$, while the same factor in §6 can be traced back, via Witten:1988hf , to an eigenvalue of a Wilson loop operator acting on ground states of $\text{SU}(2)$ Chern-Simons theory on $T^{2}$. We therefore see that the connection hinges on the observation by Vafa and Witten Vafa:1994tf about the connection between the partition function of $\mathcal{N}=4$ Super-Yang-Mills theory on $\mathbb{R}^{4}/\Gamma$ and characters of an affine Lie algebra associated to the Dynkin diagram of $\Gamma$, which in turn was obtained through earlier results of Nakajima Nakajima:1993 ; Nakajima:1994 . In particular, the number of ground states of Chern-Simons theory on $T^{2}$ is related to the number of representations of $\Gamma$. (See also a recent proof by Kojima and Tachikawa Kojima:2025hzr of the equality of the number of embeddings of $\Gamma$ in $\text{SU}(n)$ and in its Langlands dual group $\text{SU}(n)/\mathbb{Z}_{n}$, as well as other Langlands dual pairs.)
In the process of matching the results that were obtained in §5 via analytic torsion with the results that were obtained in §6 Chern-Simons theory, we were led to assume that a certain sector of the low-energy $\text{U}(1)$ Chern-Simons theory with nontrivial ‘t Hooft magnetic flux should be counted with a ()(-) sign in the total Witten index [see (157)]. We do not have an independent argument for this predicted effect. Within 5d SYM, there is no way to continuously deform the fields of the sector with ‘t Hooft magnetic flux ${\mathfrak{m}}=1$ on $S^{2}$ to the sector without it, and there is no way to continuously deform the fields of the ${\nu}=0$ [for which the solutions of (106) lead to a $\text{U}(1)$ gauge group] to the fields of ${\nu}=1$, and so the explanation for this effect seems to originate in 6d.
@@
Looking ahead, our analysis may be naturally extended in several directions. In particular, one could study the smooth orbifolds $X_{5}=S^{5}/\Gamma$ for the non-abelian $\Gamma$’s that appear in (8). Such an extension is expected to involve similar techniques, namely the computation of the Ray-Singer torsion in the background of flat connections associated to representations of $\Gamma$. Our results can also be examined in the context of 4d $\mathcal{N}=2$ $\text{Class}\;\mathcal{S}$ theories Gaiotto:2009we ; Gaiotto:2009hg , by first compactifying on $S^{2}$ (with a topological twist), leading to a gapped 4d theory. In that case, the Witten index gives information about the nontrivial topological nature of this gapped system.
@@
We have worked with an SU(2)(2) gauge group, because for SU(N)(N) with $N\geq 3$ the fermionic 0-forms will always have zero modes. This is because there are no irreducible representations of $\text{Dic}_{k}$ of dimension higher than $2$, and therefore for any representation $\mathbf{R}$ of $\text{Dic}_{k}$ there are nontrivial elements of the Lie algebra $\mathfrak{su}(N)$ (in which the 0-form takes values) that are invariant under the adjoint action of $\text{Dic}_{k}$ (embedded in SU(N)(N) via $\mathbf{R}$).
As another possible extension, one may consider more general surface operators Drukker:2020bes . It would be interesting to find explicit operators that have nontrivial matrix elements between different ground states, and a natural candidate would be a surface operator wrapped on $S^{2}$. It would also be interesting to explore the effect of the braiding group on the Wilson loops of the effective 3d theory, although the structure might not be stable under coupling constant deformation, and it might not be visible in the dual picture with the analytic torsion. In general, such studies could reveal a richer structure in the spectrum of BPS states and their associated indices.
@@ Acknowledgements.
We wish to thank Sergey Cherkis, Vi Hong, Petr Hořava and Jesus Sanchez Jr. for helpful discussions. This work has been supported by the Leinweber Institute for Theoretical Physics at UC Berkeley.
Appendix A Chern-Simons Level Shift from Integration of Massive Gluinos
In §6.5 we noted that the level of the effective $\text{SU}(2)$ Chern-Simons theory that we obtained in 3d, for the case ${\Omega}=\pm i{\mathbf{\sigma}}_{2}$, is shifted from its classical value of $k$ to (k2)(k-2) because of a mismatch between positive and negative fermionic Kaluza-Klein masses. We will show this explicitly below, and in the process, we will also show that there are no massless fermionic zero modes.
Luckily, for ${\Omega}=\pm i{\mathbf{\sigma}}_{2}$ we do not actually have to solve the full equations of motion obtained from the fermionic action (138). This is because in that case the boundary conditions (107), i.e., ${\Omega}^{-1}{\Phi}(p){\Omega}=\pm{\Phi}(\tilde{p})^{t}$ [where we wrote $p$ and $\tilde{p}$} for the antipodal points $(\frac{\pi}{2},\phi+\pi)$ and $(\frac{\pi}{2},\phi)$ in (107)] combined with the fact that ${\Omega}^{-1}\Phi(p){\Omega}=-\Phi(p)^{t}$ for every $\mathfrak{su}(2)$ Lie-algebra valued matrix, implies that the SU(2)(2) indices of fields decouple from the geometry. There is only an overall $\pm$ sign in the boundary conditions:
[…]
${\tilde{\psi}}_{{\varphi}}(p)$ $={\tilde{\psi}}_{{\varphi}}({\tilde{p}}),\quad{\tilde{\psi}}_{\theta}(p)={\tilde{\psi}}_{\theta}({\tilde{p}}),\qquad(i=1,2,{\phi}),$
and
[…]
${\tilde{\psi}}_{i}(p)$ $=-{\tilde{\psi}}_{i}({\tilde{p}}),\quad{\tilde{\rho}}(p)=-{\tilde{\rho}}({\tilde{p}}),\qquad(i,j=1,2,{\phi}).$
Thus, we can lift the problem back to the Dirac equation on $S^{3}/\mathbb{Z}_{2k}$, but with the extra condition that the modes are odd or even (depending on the $\mp$ sign) under the $\mathbb{Z}_{2}$ isometry that flips the orientation of the fiber. We are therefore back to the Dirac-Kähler equations (134)-(136) on $\mathbb{R}\times S^{2}\times S^{3}/\mathbb{Z}_{2k}$, with the extra conditions that the fields ${\tilde{\rho}},{\tilde{\psi}}_{\mu},{\tilde{\chi}}_{\mu\nu}$ are
  1. 1.
    invariant under translations in the direction of the fiber, i.e., along the Killing vector $\partial/\partial{\varphi}$, and
  2. 2.
    odd under the discrete isometry given by the action (103) of $s$ on $S^{3}$.
We need to find the Kaluza-Klein modes of the reduction of such a twisted spinor from 6d to a low-energy 3d twisted spinor on $\mathbb{R}\times S^{2}$. Firstly, we note that a 3d spinor has two components, and if it is also a doublet of an SU(2)(2) R-symmetry, we can twist it by identifying the R-symmetry with the holonomy group of $S^{2}$, thus converting the spinor into two scalars and a vector along $S^{2}$. Note that in this formalism full covariance along $\mathbb{R}\times S^{2}$ is lost, and only covariance along the directions of $S^{2}$ is maintained. With i,j{{i}},{{j}} labeling directions along $S^{2}$, after this twist, the Kaluza-Klein modes of ${\tilde{\rho}},\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$ and ${\tilde{\psi}}_{{i}}$ form a multiplet. Similarly, with i,j{{i}},{{j}} still labeling directions along $S^{2}$ and ${\mathbf{l}},{\mathbf{m}},{\mathbf{n}}$ labeling directions along $S^{3}/\mathbb{Z}_{2k}$, we find that Kaluza-Klein modes of $\psi_{\mathbf{l}},{\varepsilon_{\mathbf{l}}}^{{\mathbf{m}}{\mathbf{n}}}{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}$ and ${\tilde{\chi}}_{{{i}}{\mathbf{l}}}$ form a multiplet. This grouping into multiplets will become explicit in (170)-(175).
To describe the Kaluza-Klein reduction more concretely, we look for modes of the fields that factorize as products of a (bosonic) field on $S^{3}/\mathbb{Z}_{2k}$ and a (fermionic) field on $\mathbb{R}\times S^{2}$. Let ${\mathbf{y}}^{\mathbf{m}}$ be coordinates on $S^{3}/\mathbb{Z}_{2k}$ and let ${\mathbf{x}}^{{i}}$ be coordinates on $\mathbb{R}\times S^{2}$. We find that modes with a given Kaluza-Klein mass $\lambda$ that factorize come in two species:
  1. 1.
    modes for which the nonzero fields are proportional to components of solutions of $\star d{\mathfrak{P}}=\lambda{\mathfrak{P}}$ on $S^{3}/\mathbb{Z}_{2k}$, where $\mathfrak{P}$} is a 1-form, and
  2. 2.
    modes for which the nonzero fields are proportional to a solution $\tilde{\rho}$} of $d\star d\rho=-\lambda^{2}\star\rho$ on $S^{3}/\mathbb{Z}_{2k}$, or one of its derivatives $\partial_{\mathbf{m}}{\tilde{\rho}}$, where $\rho$ is a 0-form.
For both kinds of modes, we can first rewrite (134)-(136) by separating the derivatives along $\mathbb{R}\times S^{2}$ (i.e., with respect to ${\mathbf{x}}_{0}$ and ${\mathbf{x}}^{{i}}$) and along $S^{3}/\mathbb{Z}_{2k}$ (i.e., with respect to ${\mathbf{y}}^{\mathbf{m}}$). For factorization, we assume that the parts of the equations with derivatives along $S^{3}/\mathbb{Z}_{2k}$ satisfy the following eigenvalue conditions:
$\lambda\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$ $=$ $\nabla^{\mathbf{m}}{\tilde{\psi}}_{\mathbf{m}}\,$, (164)
$\lambda\varepsilon_{{{i}}{{j}}}{\tilde{\psi}}^{{j}}$ $=$ $2\nabla^{\mathbf{m}}{\tilde{\chi}}_{{\mathbf{m}}{{i}}}\,$, (165)
$\lambda\varepsilon_{{\mathbf{m}}{\mathbf{n}}{\mathbf{l}}}{\tilde{\chi}}^{{\mathbf{n}}{\mathbf{l}}}$ $=$ $-\partial_{\mathbf{m}}{\tilde{\rho}}-2\nabla^{\mathbf{l}}{\tilde{\chi}}_{{\mathbf{l}}{\mathbf{m}}}\,$, (166)
$\lambda{\tilde{\rho}}$ $=$ $-\varepsilon^{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\partial_{\mathbf{l}}{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}\,$, (167)
$\lambda\varepsilon_{{{i}}{{j}}}{{\tilde{\chi}}_{\mathbf{m}}}{}^{{{j}}}$ $=$ $\varepsilon_{{{i}}{{j}}}\varepsilon_{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\nabla^{\mathbf{l}}{\tilde{\chi}}^{{\mathbf{n}}{{j}}}-\frac{1}{2}\partial_{\mathbf{m}}{\tilde{\psi}}_{{i}}\,$, (168)
$\lambda{\tilde{\psi}}^{\mathbf{l}}$ $=$ $-\frac{1}{2}\varepsilon^{{{i}}{{j}}}\partial^{\mathbf{l}}{\tilde{\chi}}_{{{i}}{{j}}}-\varepsilon^{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\partial_{\mathbf{m}}{\tilde{\psi}}_{\mathbf{n}}\,.$ (169)
Then, (134)-(136) can be rewritten along $\mathbb{R}\times S^{2}$ as the following massive twisted 3d Dirac equations:
$\lambda\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$ $=$ $-\partial_{0}{\tilde{\rho}}-\partial^{{i}}{\tilde{\psi}}_{{i}}\,$, (170)
$\lambda\varepsilon_{{{i}}{{j}}}{\tilde{\psi}}^{{j}}$ $=$ $-\partial_{0}{\tilde{\psi}}_{{i}}-2\partial^{{j}}{\tilde{\chi}}_{{{j}}{{i}}}\,$, (171)
$\lambda\varepsilon_{{\mathbf{m}}{\mathbf{n}}{\mathbf{l}}}{\tilde{\chi}}^{{\mathbf{n}}{\mathbf{l}}}$ $=$ $-\partial_{0}{\tilde{\psi}}_{\mathbf{m}}-2\partial^{{i}}{\tilde{\chi}}_{{{i}}{\mathbf{m}}}\,$, (172)
$\lambda\varepsilon_{{{i}}{{j}}}{\tilde{\rho}}$ $=$ $-\partial_{0}{\tilde{\chi}}_{{{i}}{{j}}}-\frac{1}{2}\varepsilon^{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\varepsilon_{{{k}}{{j}}}\partial^{{k}}{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}\,$, (173)
$\lambda\varepsilon_{{{i}}{{j}}}{{\tilde{\chi}}_{\mathbf{m}}}{}^{{{j}}}$ $=$ $-\partial_{0}{\tilde{\chi}}_{{\mathbf{m}}{{i}}}+\varepsilon_{{{i}}{{k}}}\varepsilon_{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\partial^{{k}}{\tilde{\chi}}^{{\mathbf{n}}{\mathbf{l}}}+\frac{1}{2}\partial_{{i}}{\tilde{\psi}}_{\mathbf{m}}\,$, (174)
$\lambda{\tilde{\psi}}^{\mathbf{l}}$ $=$ $-\varepsilon^{{\mathbf{l}}{\mathbf{m}}{\mathbf{n}}}\partial_{0}{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}-\frac{1}{2}\varepsilon_{{{i}}{{k}}}\partial^{{k}}{\tilde{\chi}}^{{{i}}{\mathbf{l}}}\,.$ (175)
The 6 equations (164)-(169) are actually 3 decoupled pairs: (164) and (169) only involve ${\tilde{\chi}}_{{{i}}{{j}}}$ and ${\tilde{\psi}}_{\mathbf{m}}$, (165) and (168) only involve ${\tilde{\chi}}_{{\mathbf{m}}{{i}}}$ and ${\tilde{\psi}}_{{i}}$, and (166) and (167) only involve $\tilde{\rho}$} and ${\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}$. We will begin by analyzing the solutions to the first pair, (164) and (169).
For a fixed point on $S^{2}\times\mathbb{R}$ (labeled by ${\mathbf{x}}^{{i}}$), if we take ${\tilde{\psi}}_{\mathbf{m}}d{\mathbf{y}}^{\mathbf{m}}$ to be proportional to an eigenform $\mathfrak{P}$} of $\star d$, satisfying $\star d{\mathfrak{P}}=-\lambda{\mathfrak{P}}$, and $\lambda\neq 0$, then $\partial^{\mathbf{m}}{\tilde{\psi}}_{\mathbf{m}}$ in (164) vanishes, since it is proportional to $d\star(\star d{\mathfrak{P}})=0$. (169) will then immediately follow from $\star d{\mathfrak{P}}=\lambda{\mathfrak{P}}$. We can set all other fields to zero, and such mode will therefore have
${\tilde{\rho}}=0,\quad{\tilde{\chi}}_{{{i}}{{j}}}=0,\quad{\tilde{\chi}}_{{\mathbf{m}}{{i}}}=0,\quad{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}=0,\quad{\tilde{\psi}}_{{i}}=0,\quad{\tilde{\psi}}_{\mathbf{m}}\sim{\mathfrak{P}}_{\mathbf{m}}.$
On the other hand, if $\lambda=0$, and $\star d{\mathfrak{P}}=-\lambda{\mathfrak{P}}=0$, then $\mathfrak{P}$} is an exact 1-form, since $H^{1}(S^{3}/\mathbb{Z}_{2k},\mathbb{R})=0$. Then (164) shows that $\mathfrak{P}$} is also coexact, and hence vanishes. (169) shows that $\varepsilon^{{{i}}{{j}}}\partial^{\mathbf{l}}{\tilde{\chi}}_{{{i}}{{j}}}=0$, and hence $\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$ is a constant on $S^{3}/\mathbb{Z}_{2k}$ (at fixed ${\mathbf{x}}^{{i}}$). Thanks to the additional conditions (163), we find that $\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}=0$, and hence ${\tilde{\chi}}_{{{i}}{{j}}}$ must vanish as well. Thus there are no modes of this kind with $\lambda=0$, i.e., no zero modes, as argued in §3.
Now assume that $\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}\neq 0$ and set ${{\upsilon}}=\frac{1}{2}\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$. (We are still at a fixed point on $S^{2}\times\mathbb{R}$.) Then, (169) becomes:
$\lambda{\mathfrak{P}}=d{{\upsilon}}+\star d{\mathfrak{P}},$ (176)
and (164) becomes
$\lambda{{\upsilon}}=\star d\star{\mathfrak{P}}.$ (177)
Plugging (176) into (177) we find $\lambda^{2}{{\upsilon}}=\star d\star d{{\upsilon}}$, and so $\upsilon$}} is an eigenfunction of the Laplacian on $S^{3}/\mathbb{Z}_{2k}$ with eigenvalue $\lambda^{2}$. Because we also need to require $\upsilon$}} to be invariant under translations along $\partial/\partial{\varphi}$, i.e., along the Hopf fiber, we see that $\upsilon$}} is the pullback to $S^{3}/\mathbb{Z}_{2k}$ of a spherical harmonic $Y_{lm}(\theta,\phi)$ on the base of the Hopf fibration, and therefore $\lambda=\pm\sqrt{l(l+1)}$. Moreover, by (163) only odd $l$’s are allowed, because if $p$ is parameterized by $(\theta,\phi)$ then $\tilde{p}$} is parameterized by $(\pi-\theta,\phi+\pi)$, and $Y_{lm}(\pi-\theta,\phi+\pi)=(-1)^{l}Y_{lm}(\theta,\phi)$. Thus, there are no zero modes, and the $\lambda$’s come in pairs of positive and negative masses. From (176) we get $\lambda\star d{\mathfrak{P}}=\star d\star d{\mathfrak{P}}$, and plugging this back to the RHS of (176), and using (177) for $\upsilon$}} we find
$\lambda^{2}{\mathfrak{P}}=\lambda d{{\upsilon}}+\star d\star d{\mathfrak{P}}=d\star d\star{\mathfrak{P}}+\star d\star d{\mathfrak{P}}=\triangle_{1}{\mathfrak{P}},$
where $\triangle_{1}=d\star d\star+\star d\star d$.
The Hodge decomposition of $\mathfrak{P}$} into exact and coexact 1-forms, which incidentally is given by (176) for $\lambda\neq 0$, separates $\mathfrak{P}$} into exact and coexact eigenforms of $\triangle_{1}$, and applying $\star d$ to (176) we get $\lambda(\star d{\mathfrak{P}})=\star d\star d{\mathfrak{P}}$, which shows that $\star d{\mathfrak{P}}$ is an eigenform of $\star d$ with eigenvalue $\lambda$. But, as we will soon see below, the eigenvalues of $\star d$ are integers while $\pm\sqrt{l(l+1)}$ is never an integer for integer $l\geq 1$. Therefore, we must have $\star d{\mathfrak{P}}=0$ or $\lambda=0$. If $\star d{\mathfrak{P}}=0$ and $\lambda\neq 0$, then from (176) again we find ${\mathfrak{P}}=-(1/\lambda)d{{\upsilon}}$ and so the solution to (164) and (169) with eigenvalue $\lambda=\pm\sqrt{l(l+1)}$ and $l\geq 1$ takes the form
${\tilde{\rho}}$ $=0,\quad{\tilde{\chi}}_{{{i}}{{j}}}=\epsilon_{{{i}}{{j}}}{{\upsilon}},\quad{\tilde{\chi}}_{{\mathbf{m}}{{i}}}=0,\quad{\tilde{\chi}}_{{\mathbf{m}}{\mathbf{n}}}=0,\quad{\tilde{\psi}}_{{i}}=0,$
${\tilde{\psi}}_{\mathbf{m}}$ $=-\frac{1}{\lambda}\partial_{\mathbf{m}}{{\upsilon}},\quad{{\upsilon}}\propto Y_{lm}(\theta,\phi),\quad\lambda=\pm\sqrt{l(l+1)}$ (178)
If $\lambda=0$ then (176) and (177) imply that $\upsilon$}} is harmonic, and hence constant, and it follows that $\mathfrak{P}$} is both exact and coexact, and hence vanishes. We also saw above that if $\upsilon$}} is constant, it must vanish by (163).
We will now show that the eigenvalues $\lambda$ of $\lambda{\mathfrak{P}}=\star d{\mathfrak{P}}$ are odd integers. We set
${\mathfrak{P}}=f(\theta,\phi)(d\chi+2k\sin^{2}\left(\tfrac{\theta}{2}\right)d\phi)+{\mathcal{W}}_{\theta}(\theta,\phi)d\theta+{\mathcal{W}}_{\phi}(\theta,\phi)d\phi.$
which enforces invariance under translations along fiber $\partial/\partial\chi$. Here $f$ is (the pullback to $S^{3}/\mathbb{Z}_{2k}$ of) a globally defined 0-form on the base $S^{2}$ of the Hopf fibration $S^{3}/\mathbb{Z}_{2k}\rightarrow S^{2}$, and ${\mathcal{W}}={\mathcal{W}}_{\theta}d\theta+{\mathcal{W}}_{\phi}d\phi$ is (the pullback to $S^{3}/\mathbb{Z}_{2k}$ of) a globally defined $1$-form on the same $S^{2}$. We now need to solve
$\star df=\lambda{\mathcal{W}},\qquad\star d{\mathcal{W}}=-(\lambda+1)f,$ (179)
where the $\star$ in (179) is the 2d Hodge dual (along the base $S^{2}$ of the Hopf fibration). These imply $\star d\star df=-\lambda(\lambda+1)f$, whose solutions are spherical harmonics $Y_{lm}$ with $\lambda=l$ or $\lambda=-l-1$. The additional requirement of being even under $s$ is that $f(\pi-\theta,\phi+\pi)=f(\theta,\phi)$, which requires $l$ to be even. For $l>0$, the modes come in pairs of even and odd eigevalues $l>0$ and l1<0-l-1<0, and those add up to a net shift of 0 to the low-energy effective Chern-Simons level. For $l=0$ we get a constant $f$ and $d{\mathcal{W}}$ is proportional to $\sin\theta d\theta\wedge d\phi$. But $\mathcal{W}$} has to be globally defined and $\sin\theta d\theta\wedge d\phi$ is not exact, so there is no solution for $\mathcal{W}$} with $l=0$ and the mode with $\lambda=-1$ is unpaired.
The solution of (166) and (167) is similar to the solution we discussed above of (164) and (169), with $\tilde{\rho}$} playing the role of $\varepsilon^{{{i}}{{j}}}{\tilde{\chi}}_{{{i}}{{j}}}$, and $\varepsilon_{{\mathbf{m}}{\mathbf{n}}{\mathbf{l}}}{\tilde{\chi}}^{{\mathbf{n}}{\mathbf{l}}}$ playing the role of ${\tilde{\psi}}_{\mathbf{m}}$. The only remaining pair of equations to solve is therefore (165) and (168). This is again a similar pair of equations, with a vector index i{{i}} (on $S^{2}$ added), so that for i{{i}} fixed $\varepsilon_{{{i}}{{j}}}{\tilde{\psi}}^{{j}}$ now plays the role of $\varepsilon^{{{k}}{{j}}}{\tilde{\chi}}_{{{k}}{{j}}}$, and ${\tilde{\chi}}_{{\mathbf{m}}{{i}}}$ now plays the role of ${\tilde{\psi}}_{\mathbf{m}}$. Altogether, there is exactly one unpaired 3d fermion mass, leading to a shift of the Chern-Simons level to k2k-2 as claimed.